Operations Report – January 18th

Crew 219 Operations Report 18-JAN-2020

Sol: 13

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 13.5

From what time last night: 17:30

To what time this morning: 07:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 80%

10:00 = 95%; 11:00 = 93%; 12:30 = 91%; 13:45 = 90%; 15:30 = 86% 16:30 = 82%

Diesel Reading – 55%

Station Propane Reading- 29%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147887,4

Water (static tank) (gal) – 206

Static to loft Pump used –Yes

Water in Green Hab (gal): 133.39

Water in Science Dome (gal): 0

Toilet Tank Emptied –Yes

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- No

Starting Hours: 126.4

Beginning Charge: 100%

Currently Charging: Yes

Ending Hours:

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 132.4

Beginning Charge: 100%

Currently Charging: Yes

Ending Hours:

Ending Charge:

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Yes, by Outpost for to get Crew 220

CrewCar used and why, where? On campus.

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: Nominal

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: See HSO report.

Questions, concerns, and requests to Mission Support: